# Integral cohomology rings of weighted Grassmann orbifolds and rigidity properties

### 11 Mar 2024 v2

@@ 4. Torsion’s in the integral cohomology of weighted Grassmann orbifold
    ${\bf{b}}=(\alpha\beta\gamma,\alpha\beta\gamma,(\alpha\beta-\alpha+1)\gamma,\alpha\beta\gamma,\alpha\gamma,\gamma,\alpha\gamma,\gamma,\alpha\gamma,\gamma)\in(\mathbb{Z}_{\geq 1})^{10}.$
    Then the integral cohomology of $\mbox{Gr}_{\bf{b}}(2,5)$ has no torsion using Theorem 4.1.
@@
    Let $\alpha,\beta,\gamma,s$ ant $t$ be positive integers. Consider
    ${\bf{b}}=(\gamma,(1+\beta)st\alpha,(1+s\beta)t\alpha,(1+\beta)st\alpha,st\alpha,t\alpha,t\alpha,\alpha,\alpha,\alpha)\in(\mathbb{Z}_{\geq 1})^{10}.$
    Then the integral cohomology of $\mbox{Gr}_{\bf{b}}(2,5)$ has no torsion using Theorem 4.1.
  3. (3)
    Let $a$ and $c$ be two positive integer such that $\gcd\{a,c\}=1$. Let $s=a+c$. For any positive integer $\alpha$ such that $s-\alpha$ divides $\alpha$ and $c$ divides $\alpha$ then the $\mbox{Gr}_{\bf{b}}(2,5)$ has no torsion where ${\bf{b}}=(a,\alpha,\alpha,\alpha,s-\alpha,s-\alpha,s-\alpha,c,c,c)$. For example, ${\bf{b}}=(5,4,4,4,2,2,2,1,1,1)$, ${\bf{b}}=(28,24,24,24,6,6,6,2,2,2)$
Now we define divisive weighted Grassmann orbifold in terms of Plücker weight vector and Plücker permutation following [7].
@@
Let $\alpha$ and $\beta$ be two positive integers. Consider a Plücker weight vector ${\bf b}=(\alpha,\alpha\beta,\alpha\beta,\alpha,\alpha,\alpha\beta)$. Define a Plücker permutation $\sigma$ on $\{0,1,\dots,5\}$ by $\sigma(0)=5$, $\sigma(5)=0$ and $\sigma(i)=i$ for $i=1,2,3,4$. Then $\mbox{Gr}_{\bf b}(2,4)$ is a divisive weighted Grassmann orbifold.
The proof of the following theorem follows from [7, Theorem 3.19].
@@
A divisive weighted Grassmann orbifold has a CW complex structure with even dimensional cells $\{\mathbb{C}^{d(\lambda^{i})}\}_{i=0}^{m}$.
@@
In this section, first we revisit the $T^{n}$-equivariant cohomology ring of divisive weighted Grassmann orbifolds with integer coefficients as introduced in [7]. Our focus lies on studying the multiplication formulae in the integral equivariant cohomology ring of a divisive weighted Grassmann orbifold. We consider a basis of the aforementioned ring as a $\mathbb{Z}[y_{1},\dots,y_{n}]$ module and compute all the equivariant structure constants concerning that basis. Moreover, we demonstrate the positivity of the equivariant structur constants.
Let $T^{n}=(S^{1})^{n}$ be the $n$-dimensional compact abelian torus. Consider the $T^{n}$-action on $\mbox{Gr}_{\bf{b}}(k,n)$ as the restriction of the (*$(\mathbb{C}^{*})^{n}$-action defined in (3.1). Let $\mbox{Gr}_{\bf b}(k,n)$ be a divisive weighted Grassmann orbifold. We can assume that $b_{i}$ divides $b_{i-1}$ for all $i=1,2,\dots,m$ using Theorem 3.4. Define $d_{ij}:=\frac{b_{i}}{b_{j}}$ for $i<j$ and recall $R(\lambda^{i})$ from (2.11).
@@
[…]
Here $\mathbb{Z}[y_{1},\dots,y_{n}]=H^{*}_{T^{n}}(\{pt\};\mathbb{Z})$ and $Y_{\lambda}:=\sum_{i=1}^{k}y_{\lambda_{i}}$ for the Schubert symbol $\lambda=(\lambda_{1},\ldots,\lambda_{k})$.
For the particular case, ${\bf{b}}=(1,1,\dots,1)$ the $T^{n}$-equivariant cohomology ring of $\mbox{Gr}_{\bf b}(k,n)$ coincide with the $T^{n}$-equivariant cohomology ring of $\mbox{Gr}(k,n)$ considered in [14]. In [14, Section 2], the authors computed the multiplication formulae for the integral equivariant cohomology ring $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$ of the Grassmann manifold $\mbox{Gr}(k,n)$.
5.1. Knutson-Tao’s results on $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$
Corresponding to every Schubert symbol $\lambda=(\lambda_{1},\dots,\lambda_{k})$, one can consider a string consisting of $k$ ones and $(n-k)$ zeroes in the following order. One for each $\lambda_{i}$-th position and zero for the remaining $(n-k)$ positions. In this way, we can identify the set of all Schubert symbols $\lambda=(\lambda_{1},\dots,\lambda_{k})$ and the set of strings consisting of $k$ ones and $(n-k)$ zeroes in arbitrary order, as explored by Knutson and Tao, see [14]. We recall an inversion of a Schubert symbol $\lambda$ is a pair $(s,s^{\prime})$ such that $s\in\lambda,s^{\prime}\notin\lambda$ and $s<s^{\prime}$. We denote the set of all inversions of $\lambda$ by $\rm{inv}(\lambda)$. If $(s,s^{\prime})\in\rm{inv}(\lambda)$, then $(s,s^{\prime})\lambda$ denote the Schubert symbol obtained by replacing $s$ by $s^{\prime}$ in $\lambda$ and ordering the later set.
For every $i\in\{0,1,\dots,m\}$ consider a subset of Schubert symbols
$I(\lambda^{i})=\{\lambda^{j}:\lambda^{j}=(s,s^{\prime})\lambda^{i}\text{ for }(s,s^{\prime})\in\rm{inv}(\lambda^{i})\}.$
Let $d^{\prime}(\lambda^{i})$ denotes the cardinality of the set $I(\lambda^{i})$. Then $d(\lambda^{i})+d^{\prime}(\lambda^{i})=k(n-k)$.
Proposition 5.2.
[14] There exist an equivariant Schubert basis $\{\widetilde{S}_{\lambda^{i}}\}_{i=0}^{m}$ for the equivariant cohomology ring $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$. Each basis element $\widetilde{S}_{\lambda^{i}}$ is uniquely determined the following conditions
  1. (1)
    $\widetilde{S}_{\lambda^{i}}|_{\lambda^{j}}\neq 0\implies\lambda^{j}\preceq\lambda^{i}$.
  2. (2)
    $\widetilde{S}_{\lambda^{i}}|_{\lambda^{i}}=\prod_{\lambda^{\ell}\in I(\lambda^{i})}(Y_{\lambda^{\ell}}-Y_{\lambda^{i}})$.
  3. (3)
    $\widetilde{S}_{\lambda^{i}}|_{\lambda^{j}}$ is a homogeneous polynomial of $y_{1},y_{2},\dots,y_{n}$ of degree $d^{\prime}({\lambda^{i}})$.
We have the following multiplication rule in the equivariant cohomology ring $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$
(5.1) $\widetilde{S}_{\lambda^{i}}\widetilde{S}_{\lambda^{j}}=\sum_{\lambda^{\ell}}\widetilde{C}_{i~{}j}^{\ell}\widetilde{S}_{\lambda^{\ell}}.$
In their study [14], Knutson and Tao conducted explicit computations of the equivariant structure constant $\widetilde{C}_{i~{}j}^{\ell}$ with respect to the basis $\{\widetilde{S}_{\lambda^{i}}\}$ for $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$. We refer to [14, Section 1] for the definition of puzzle pieces and equivariant puzzle pieces made by triangles and rhombus with all edges labeled 0 or 1. One can use these pieces to construct a puzzle $P$ as a decomposition of a labeled equilateral triangle into puzzle pieces and equivariant puzzle pieces. By a labeled equilateral triangle, we mean an upward-pointing equilateral triangle with integer side length $n$ and each side is labeled by a string consisting of $k$ ones and $(n-k)$ zeroes. The labeled equilateral triangle induced from the puzzle $P$ is denoted by $\partial P$, the boundary of $P$. Let $P$ be a puzzle with $\partial P=\Delta_{\lambda^{i}~{}\lambda^{j}}^{\lambda^{\ell}}$. Using [14] the number of equivariant puzzle pieces in $P$ is $|P|:=d^{\prime}(\lambda^{i})+d^{\prime}(\lambda^{j})-d^{\prime}(\lambda^{\ell})\geq 0$. Let $p_{s}$ be equivariant puzzle pieces of $P$, for $s=1,2,\dots,|P|.$ Then $wt(P)\colon=\prod_{s=1}^{|P|}wt(p_{s})$, where $wt(p_{s})=y_{a}-y_{c}$ for $a>c$ and $wt(P)=1$ if $|P|=0$.
Proposition 5.3.
[14, Theorem 2] The equivariant structure constant is equal to the sum of the weights of all puzzle $P$ such that $\partial P=\Delta_{\lambda^{i}~{}\lambda^{j}}^{\lambda^{\ell}}$.
$\widetilde{C}_{i~{}j}^{\ell}=\sum_{P~{}:~{}\partial P=\Delta_{\lambda^{i}~{}\lambda^{j}}^{\lambda^{\ell}}}wt(P)$
5.2. Structure constants for the equivariant cohomology of $\mbox{Gr}_{\bf b}(k,n)$
In this subsection, we compute the multiplication structure in the equivariant cohomology ring $H_{T^{n}}^{*}(\mbox{Gr}_{\bf b}(k,n);\mathbb{Z})$.
Proposition 5.4.
[7] There exist an equivariant Schubert basis $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ for the equivariant cohomology ring $H_{T^{n}}^{*}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Z})$ of a divisive weighted Grassmann orbifold. Moreover, each basis element ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}$ is uniquely determined by the following conditions.
@@ 5.2. Structure constants for the equivariant cohomology of $\mbox{Gr}_{\bf b}(k,n)$
Our goal is to compute the equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ explicitly. The author and Sarkar have studied some interesting property of ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ in [7, Lemma 5.5].
For the equivariant cohomology ring $H^{*}_{T^{n}}(\mbox{Gr}(k,n);\mathbb{Z})$ we denote the basis $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ by $\{\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$. Let $\sigma_{r}\in S_{n}$ be the permutation defined by
$\sigma_{r}:=\begin{pmatrix}1&2&3&\cdots&n-1&n\\ n&n-1&n-2&\cdots&2&1\end{pmatrix}.$
In other words, $\sigma_{r}(i)=n+1-i$. Note that $\sigma_{r}^{2}(i)=i$. Then $\sigma_{r}$ induces a permutation on $\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$ defined by $\sigma_{r}y_{i}=y_{n+1-i}$. Thus $\sigma_{r}$ induces a permutation on $H_{T^{n}}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Z})$.
The following proposition gives the relation between two bases of $H_{T^{n}}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Z})$.
Proposition 5.5.
$\widetilde{\mathbb{S}}_{\sigma_{r}\lambda^{i}}|_{\sigma_{r}\lambda^{j}}=\sigma_{r}(\widetilde{S}_{\lambda^{i}}|_{\lambda^{j}}).$
Proof.
Note that $\lambda^{\ell}\in R(\sigma_{r}\lambda^{i})$ if and only if $\sigma_{r}\lambda^{\ell}\in I(\lambda^{i})$. Thus
$\widetilde{\mathbb{S}}_{\sigma_{r}\lambda^{i}}|_{\sigma_{r}\lambda^{i}}$ $=\prod_{\lambda^{\ell}\in R(\sigma_{r}\lambda^{i})}(Y_{\lambda^{\ell}}-Y_{\sigma_{r}\lambda^{i}})$
$=\prod_{\sigma_{r}\lambda^{\ell}\in I(\lambda^{i})}(Y_{\sigma_{r}(\sigma_{r}\lambda^{\ell})}-Y_{\sigma_{r}\lambda^{i}})$
$=\prod_{\lambda^{s}\in I(\lambda^{i})}(Y_{\sigma_{r}\lambda^{s}}-Y_{\sigma_{r}\lambda^{i}})=\sigma_{r}(\widetilde{S}_{\lambda^{i}}|_{\lambda^{i}}).$
Also, both are of degree $d(\sigma_{r}\lambda^{i})=d^{\prime}(\lambda^{i})$. Now the proof follows from the uniqueness of the equivariant Schubert basis $\widetilde{\mathbb{S}}_{\lambda^{i}}$. ∎
@@
$\widetilde{S}_{\lambda^{i}}$ represents the equivariant cohomology class in $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$ corresponding to the closure of the cell $E(\lambda^{i})$, see [14, Section 2.3]. Using Proposition 5.5, $\widetilde{\mathbb{S}}_{\lambda^{i}}|_{\lambda^{j}}=\sigma_{r}(\widetilde{S}_{\sigma_{r}\lambda^{i}}|_{\sigma_{r}\lambda^{j}}).$ Thus one can think $\widetilde{\mathbb{S}}_{\lambda^{i}}$ as the equivariant cohomology class in $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$ corresponding to the closure of the cell $\sigma_{r}E(\sigma_{r}\lambda^{i})$.
Let $\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ be the equivariant structure constant corresponding to the basis $\{\widetilde{\mathbb{S}}_{\lambda^{i}}\}$. Then we have the following formulae for $\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$. Let $p_{s}$ be an equivariant puzzle piece such that $wt(p_{s})=y_{a}-y_{c}$ for some $a>c$. Define $\sigma_{r}wt(p_{s}):=y_{n+1-a}-y_{n+1-c}$ for $n+1-a<n+1-c$ and $\sigma_{r}wt(P):=\prod_{s=1}^{|P|}\sigma_{r}wt(p_{s})$.
Lemma 5.7.
$\widetilde{\mathscr{C}}_{i~{}j}^{\ell}=\sum_{P~{}:~{}\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{\ell}}}\sigma_{r}wt(P).$
Proof.
Apply $\sigma_{r}$ from left in (5.1) we get
$(\sigma_{r}\widetilde{S}_{\lambda^{i}})(\sigma_{r}\widetilde{S}_{\lambda^{j}})$ $=\sum_{\lambda^{\ell}}\sigma_{r}\widetilde{C}_{i~{}j}^{\ell}(\sigma_{r}\widetilde{S}_{\lambda^{\ell}})$
$(\widetilde{\mathbb{S}}_{\sigma_{r}\lambda^{i}})(\widetilde{\mathbb{S}}_{\sigma_{r}\lambda^{j}})$ $=\sum_{\lambda^{\ell}}\sigma_{r}\widetilde{C}_{i~{}j}^{\ell}(\widetilde{\mathbb{S}}_{\sigma_{r}\lambda^{\ell}})$
${\mathscr{\widetilde{C}}}_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{\ell}}$ $=\sigma_{r}{\widetilde{C}}_{i~{}j}^{\ell}.$
Now the proof follows using Proposition 5.3. ∎
Next we use Lemma 5.7 to calculate the equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ for any arbitrary Plücker weight vector ${\bf{b}}$.
Recall (*$(\mathbb{C}^{*})^{n+1}$ action on $Pl(k,n)$ from (2.6) which is given by the map
$\rho\colon(\mathbb{C}^{*})^{n+1}\to(\mathbb{C}^{*})^{m+1}$
defined by $\rho(t_{1},t_{2},\dots,t_{n},t)=(t\cdot t_{\lambda^{0}},t\cdot t_{\lambda^{1}},\dots,t\cdot t_{\lambda^{m}}).$
Consider a weighted Grassmann orbifold $\mbox{Gr}_{\bf b}(k,n)$. Using Proposition 2.9, there exists $W=(w_{1},\dots,w_{n})\in\mathbb{Z}^{n}$ and $a\in\mathbb{Z}_{\geq 1}$ such that $b_{i}=w_{\lambda^{i}}$. Note that
$\rho(WD)=\{(t^{b_{0}},t^{b_{1}},\dots,t^{b_{m}})~{}|~{}t\in\mathbb{C}^{*}\},$
where $WD$ is the subgroup of (*$(\mathbb{C}^{*})^{n+1}$ defined in (2.7). Thus the action of the subgroup $\rho(WD)$ on $\rho((\mathbb{C}^{*})^{n+1})$ is same as the ${\bf b}$-action of $\mathbb{C}^{*}$. Let $T^{n+1}=(S^{1})^{n+1}$ and define $T_{\bf b}:=\frac{\rho(T^{n+1})}{{\bf b}\text{-action of }S^{1}}$. Consequently there is a $T_{\bf b}$ action on $\mbox{Gr}_{\bf{b}}(k,n)$. Note that $H^{*}(BT_{{\bf{b}}};\mathbb{Q})=\mathbb{Q}[x_{1},\dots,x_{n}]$, where $x_{1}=y_{1}-\frac{w_{1}}{a}z,\dots,x_{n}=y_{n}-\frac{w_{n}}{a}z$.
Using Vitoris Begle mapping theorem, the authors in [1, Proposition 3.1] proved that there is an isomorphism between the two equivariant cohomology ring with rational coefficients
@@
as rings over the polynomial rings $H^{*}(BT^{n};\mathbb{Q})$.
The advantage of taking the $T_{\bf{b}}$ action on $\mbox{Gr}_{\bf{b}}(k,n)$ is
@@
is isomorphism as rings over the polynomial rings $H^{*}(BT_{\bf{b}};\mathbb{Q})$.
The authors in [1] computed the $T^{n+1}$-equivariant cohomology ring of $Pl(k,n)$ with rational coefficients. We denote $H^{*}(BT^{n+1};\mathbb{Q})=\mathbb{Q}[y_{1},\dots,y_{n},z]$.

### 09 Mar 2026 v3

@@ 4. Torsion’s in the integral cohomology of weighted Grassmann orbifold
    ${\bf{b}}=(\gamma,\alpha\gamma,\alpha\gamma,\alpha\beta\gamma,\gamma,\gamma,(\alpha\beta-\alpha+1)\gamma,\alpha\gamma,\alpha\beta\gamma,\alpha\beta\gamma)\in(\mathbb{Z}_{\geq 1})^{10}.$
    Then the integral cohomology of the $\mbox{Gr}_{\bf{b}}(2,5)$ has no torsion using Theorem 4.1.
@@
    Let $\alpha,\beta,\gamma,s$ ant $t$ be positive integers with $\alpha+\gamma=t\alpha(1+s+s\beta)$. Consider
    ${\bf{b}}=(\alpha,\alpha,t\alpha,(1+\beta)st\alpha,\alpha,t\alpha,(1+\beta)st\alpha,t\alpha,(1+\beta)st\alpha,\gamma)\in(\mathbb{Z}_{\geq 1})^{10}.$
    Then the integral cohomology of the $\mbox{Gr}_{\bf{b}}(2,5)$ has no torsion using Theorem 4.1.
Next, we define divisive weighted Grassmann orbifold in terms of Plücker weight vector and Plücker permutation following [7].
@@
Let $\alpha$ and $\beta$ be two positive integers. Consider a weight vector ${\bf b}=(\alpha,\alpha\beta,\alpha\beta,\alpha,\alpha,\alpha\beta)$. Then $\bf b$} is a Plücker weighte vector for $k=2,n=4$ following Example 2.5. Define a Plücker permutation $\sigma$ on $\{0,1,2,3,4,5\}$ by $\sigma(0)=5$, $\sigma(5)=0$ and $\sigma(i)=i$ for $i=1,2,3,4$. Then $\mbox{Gr}_{\bf b}(2,4)$ is a divisive weighted Grassmann orbifold.
Remark 4.8.
Consider a Plücker weight vector ${\bf b}=(1,12,12,1,1,12)$ for $k=2,n=4$ corresponding to W=(5,6,5,5)W=(5,-6,5,5) and $a=2$. Consider the scaled Plücker weight vector $7{\bf b}=(7,84,84,7,7,84)$ which corresponds to W=(41,36,41,41)W=(41,-36,41,41) and $a=2$. Now $\mbox{Gr}_{\bf b}(2,4)$ isomorphic to $\mbox{Gr}_{7\bf b}(2,4)$ by Lemma 3.7. Using Lemma 3.4, $\mbox{Gr}_{7\bf b}(2,4)$ isomorphic to $\mbox{Gr}_{\bf b^{\prime}}(2,4)$ where ${\bf b^{\prime}}=(84,84,84,7,7,7)$ that correspond to $W=(80,3,3,3)$ and $a=1$. All the weighted Grassmann orbifolds appearing above are divisive.
The proof of the following theorem follows from same argument as in [7, Theorem 3.19].
@@
A divisive weighted Grassmann orbifold has a CW complex structure with even dimensional cells $\{\mathbb{C}^{d^{\prime}(\lambda^{i})}\}_{i=0}^{m}$.
Proof.
From (2.10), we see that if $z]=[(z_{0},\dots,z_{m})]\in E(\lambda^{i}),$ then $z_{j}=0$ for every $j<i$. The $G(b_{i})$-action on $E(\lambda^{i})$, described in (2.12) becomes trivial whenever $b_{i}$ divides $b_{j}$ for all $\lambda^{j}\in I(\lambda^{i})$. Let $\mbox{Gr}_{\bf b}(k,n)$ is a divisive weighted Grassmann orbifold. Consider the $q$-CW complex structure on $\mbox{Gr}_{\sigma\bf b}(k,n)$ given by $\{\frac{E(\lambda^{i})}{G(b_{\sigma(i)})}\}_{i=0}^{m}$ following Remark 3.5. Since $\mbox{Gr}_{\bf b}(d,n)$ is divisive, there exists a Plücker permutation $\sigma$ such that $b_{\sigma(i)}$ divides $b_{\sigma(j)}$ for every $\lambda^{j}\in I(\lambda^{i})$. Thus the action of $G(b_{\sigma(i)})$ on $E(\lambda^{i})$ is trivial. Consequently,
$\frac{E(\lambda^{i})}{G(b_{\sigma(i)})}\cong E(\lambda^{i})\cong\mathbb{C}^{d^{\prime}(\lambda^{i})},$
where $\mathbb{C}^{d^{\prime}(\lambda^{i})}$ is defined in (2.16). ∎
@@
In this section, we explore the $T^{n}$-equivariant cohomology ring of divisive weighted Grassmann orbifolds with integer coefficients. Our main focus lies on studying the multiplication formulae in the integral equivariant cohomology ring of divisive weighted Grassmann orbifolds. We consider the equivariant Schubert basis of the aforementioned ring as a $\mathbb{Z}[y_{1},\dots,y_{n}]$ algebra and compute all the structure constants concerning that basis. Moreover, we demonstrate the positivity of the equivariant structure constants.
Let $T^{n}=(S^{1})^{n}$ be the $n$-dimensional compact abelian torus. Consider the $T^{n}$-action on $\mbox{Gr}_{\bf{b}}(k,n)$ as the restriction of the $(\mathbb{C}^{*})^{n}$-action defined in (3.1). Let $\mbox{Gr}_{\bf b}(k,n)$ be a divisive weighted Grassmann orbifold. We can assume that $b_{i}$ divides $b_{i-1}$ for all $i=1,2,\dots,m$ using Lemma 3.4. Define $d_{ij}:=\frac{b_{i}}{b_{j}}$ for $i<j$. We recall $I(\lambda^{i})$ from (2.15) and define
$R(\lambda^{i}):=\{\lambda^{j}:\lambda^{i}\in I(\lambda^{j})\}.$
Then $|R(\lambda^{i})|=d(\lambda^{i})$. For every Schubert symbol $\lambda=(\lambda_{1},\ldots,\lambda_{k})$ we define $Y_{\lambda}:=\sum_{i=1}^{k}y_{\lambda_{i}}\in\mathbb{Z}[y_{1},\dots,y_{n}]=H^{*}_{T^{n}}(\{pt\};\mathbb{Z})$.
@@
$H_{T^{n}}^{*}({\rm{Gr}}_{\bf{b}}(k,n);\mathbb{Z})=\Big\{\alpha\in\bigoplus_{j=0}^{m}{\mathbb{Z}}[y_{1},\dots,y_{n}]~\big|~(Y_{\lambda^{i}}-d_{ij}Y_{\lambda^{j}})~\mbox{divides}~(\alpha_{j}-\alpha_{i})\text{ if }\lambda^{i}\in R(\lambda^{j})\Big\}.$
Next, we give a combinatorial description of the equivariant Schubert classes ${\bf b}\widetilde{\mathbb{S}}_{\lambda^{i}}\in H_{T^{n}}^{*}({\rm{Gr}}_{\bf{b}}(k,n);\mathbb{Z})$ corresponding to every Schubert symbol $\lambda^{i}$.
Definition 5.2.
[7] An element ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\in H_{T^{n}}^{*}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Z})$ is called an equivariant Schubert class corresponding to the Schubert symbol $\lambda^{i}$ if the following conditions are satisfied.
@@
The next Proposition follows from [7, Proposition 5.3] as well as [10, Proposition 4.3].
Proposition 5.3.
There exists a unique Schubert class ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\in H_{T^{n}}^{*}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Z})$ corresponding to every Schubert symbol $\lambda^{i}$. Moreover $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ forms a basis for the equivariant cohomology ring $H_{T^{n}}^{*}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Z})$ as a $\mathbb{Z}[y_{1},\dots,y_{n}]$ algebra.
@@
The author and Sarkar have studied some properties of ${\bf{b}}\widetilde{\mathscr{C}}_{i~j}^{\ell}$ in [7, Lemma 5.5]. If ${\bf b}=(1,1,\dots,1)$, we denote the basis $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ by $\{\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$, which forms a basis for the equivariant cohomology ring $H^{*}_{T^{n}}(\mbox{Gr}(k,n);\mathbb{Z})$ as a $\mathbb{Z}[y_{1},\dots,y_{n}]$ algebra. There exists $\widetilde{\mathscr{C}}_{i~j}^{\ell}\in\mathbb{Z}[y_{1},\dots,y_{n}]$ such that
(5.2) $\widetilde{\mathbb{S}}_{\lambda^{i}}\widetilde{\mathbb{S}}_{\lambda^{j}}=\sum_{\ell=0}^{m}\widetilde{\mathscr{C}}_{i~j}^{\ell}\widetilde{\mathbb{S}}_{\lambda^{\ell}}$
Knutson-Tao found a nice combinatorial formulae of the equivariant structure constants $\widetilde{\mathscr{C}}_{i~j}^{\ell}$ in terms of puzzles in [16, Section 2]. Corresponding to every Schubert symbol $\lambda=(\lambda_{1},\dots,\lambda_{k})$, one can consider a string consisting of $k$ ones and (nk)(n-k) zeroes in the following order. One for each $(n+1-\lambda_{i})$-th position and zero for the remaining (nk)(n-k) positions. In this way, we can identify the set of all Schubert symbols $\lambda=(\lambda_{1},\dots,\lambda_{k})$ and the set of strings consisting of $k$ ones and (nk)(n-k) zeroes in arbitrary order, as explored by Knutson and Tao, see [16].
Moreover, from [16, Section 1.2] it follows that the equivariant structure constants $\widetilde{\mathscr{C}}_{i~j}^{\ell}$ defined in (5.2) can be expressed as a polynomial in $\{(y_{r}-y_{r-1}):1\leq r\leq n-1\}$ such that coefficients are non-negative integers. Let $R$ be a finite collection of elements in $\{1,2,\dots,n-1\}$. Define $u_{r}:=y_{r}-y_{r+1}$ for $r\in\{1,2,\dots,n-1\}$ and $U_{R}=\prod_{r\in R}u_{r}$. Then we can rewrite the above equivariant structure coefficient by the following.
(5.3) $\widetilde{\mathscr{C}}_{i~j}^{\ell}=\sum_{|R|=d(\lambda^{i})+d(\lambda^{j})-d(\lambda^{\ell})}K(i,j,\ell,R)U_{R}$
where $K(i,j,\ell,R)\in\mathbb{Z}_{\geq 0}$ is the coefficient of $U_{R}$ in $\widetilde{\mathscr{C}}_{i~j}^{\ell}$. Thus $K(i,j,\ell,R)=0$ if there is no monomial in the expansion of $\widetilde{\mathscr{C}}_{i~j}^{\ell}$ which is constant multiple of $U_{R}$.
@@
  1. (1)
    The elements in $R$ need not be always distinct; any elements may appear finitely many times. For instance, $U_{R}$ could be $u_{3}^{2}u_{4}^{3}$ corresponding to $R=\{3,3,4,4,4\}$.
  2. (2)
    The sign convention in the equivariant structure coefficient described in (5.3) differs from the formulae appearing in [16]. This discrepancy arises from a permutation $\sigma_{r}:\{1,2,\dots,n\}\rightarrow\{1,2,\dots,n\}$ defined by $\sigma_{r}(i)=n+1-i$. The same permutation acts on the polynomial ring $\mathbb{Z}[y_{1},\dots,y_{n}]$ via $\sigma_{r}(y_{i})=y_{n+1-i}$, and therefore induces an action on $H_{T^{n}}^{*}(\mathrm{G}(k,n);\mathbb{Z})$. If $\{S_{\lambda^{i}}\}_{i=0}^{m}$ denotes the basis of $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$ used in [16], then the two conventions are related by $\widetilde{\mathbb{S}}_{\lambda^{i}}=\sigma_{r}(\widetilde{S}_{\lambda^{i}}).$
We recall $(\mathbb{C}^{*})^{n+1}$-action on $Pl(k,n)$ from (2.6) and the subgroup $WD$ of $(\mathbb{C}^{*})^{n+1}$ from (2.7). The $T^{n+1}$ action on $Pl(k,n)$ induces an $T_{\bf b}$ action on $\mbox{Gr}_{\bf{b}}(k,n)$, where $T_{\bf b}:=\frac{T^{n+1}}{WD}$. We denote $H^{*}(BT^{n+1};\mathbb{Q})=\mathbb{Q}[y_{1},\dots,y_{n},z]$ and $H^{*}(BT_{{\bf{b}}};\mathbb{Q})=\mathbb{Q}[x_{1},\dots,x_{n}]$, where $x_{r}=y_{r}-\frac{w_{r}}{a}z$ for $1\leq r\leq n$.
There is an isomorphism between the two equivariant cohomology ring
@@
as algebras over $H^{*}(BT^{n};\mathbb{Q})=\mathbb{Q}[y_{1},\dots,y_{n}]$. Moreover, The advantage of taking the $T_{\bf{b}}$ action on $\mbox{Gr}_{\bf{b}}(k,n)$ is
@@
is isomorphism as algebras over $H^{*}(BT_{\bf{b}};\mathbb{Q})=\mathbb{Q}[x_{1},\dots,x_{n}]$, see [1, Proposition 3.1]. In addition there exists an isomorphism
$\phi^{*}:H^{*}(BT_{\bf{b}};\mathbb{Q})\to H^{*}(BT^{n};\mathbb{Q})$
defined by $\phi^{*}(x_{i})=y_{i}$. Moreover, $\phi^{*}$ induce an isomorphism
$H_{T_{\bf{b}}}^{*}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Q})\cong H_{T^{n}}^{*}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Q}).$
This follows immediately from the explicit descriptions of the corresponding GKM rings given in Theorem 3.13 and [1, Theorem 4.1] respectively.
[…]
Figure 1. A diagram explaining the algebra isomorphisms